Consider the following data for Nike Inc : In 2009 it had $19.300 million in sales with a 10% growth rate in 2010, but then slows by 1% to the long-un growth ate of 5% by 2015, Nike expects EBIT to be 10% of sales, increases in net working capital requirements to be 10% of any increases equal depreciation expenses. Nike also has $2.300 million in cash, $32 million in debt, 486 million shares outstanding, a tax rate of 24%, and a weighted average cost in sales, and capital expenditures to al of capital of 10%. a Suppose you believe Nikes initial revenue growth rate will be between 7% and 1 1% with growth slowing linearly to 5% by year 2015). What ange of prices for Nike Me stock is with these forecasts? b Sup ose you be eve N kes initia venue EIT margin wil be between 9% and 1 t % of saes. What range of prices for Nike stock is consistent with these forecasts? c Suppose you believe Nikes weighted average cost of capital is between 9.5% and 12%. What range of prices for Nike stock is consistent with these forecasts? d. What range of stock prices is consistent if you vary the estimates as in parts (a), (b), and (c) simultaneously? d 11% with growth slowing in early to 5% by year 2015 What range of prices for Nike a. Suppose you believe Nikes initial revenue growth rate will be between 7% an stock is consistent with these forecasts? Round to the nearest cent.) Lowest price share:(Round to the nearest cent.) b. Suppose you The range of prices will be Highest price per share: S(Round to the nearest cent.) believe Nikes initial revenue EBIT margin will be between 9% and 11% of sales. What range of prices for Nike stock is consistent with these Enter your answer in each of the answer boxes Chapter 13 Homework (Required
Answer #1

Hi,

Please find the answers:-

a) $57.67, $64.18

b) $55.27, $69.75

Workings are:

Information Provided:-

Cash (mn) 2300
Debt (mn) 32
shares o/s (mn) 486
Tax 24%
WACC 10%

Calculating Free Cash flows for calculating the Stock price:-

Year 2009 2010 2011 2012 2013 2014 2015
Sales (mn)                    19,300            21,230          23,141                  24,992          26,741          28,346          29,763
Growth rate 10% 9% 8% 7% 6% 5%
EBIT (10%)               2,123            2,314                     2,499            2,674            2,835            2,976
Tax (24%)                  510                555                        600                642                680                714
Increase in WC (10% of increae in sales)                  193                191                        185                175                160                142
Free cash flow               1,420            1,568                     1,714            1,857            1,994            2,120
PV of cash flows               1,291            1,296                     1,288            1,269            1,238            1,197
Calculating the current stock price:
Terminal Value of the firm is equal to TV = (FCFn x (1 + g)) / (WACC – g)
in this case we have g as 5% and WACC as 10% and FCF at n as 2120 substituting these in the formula we get =2120*((1+0.05)/(0.1-0.05)) Terminal Value of Nike as: $44520 mn

Now we will compute Present value of all cash flows and terminal value

PV of Terminal Value =44520/(1.1)^6            25,130
Sum of PV of Cash flows above               7,578
Total PV of cash flows (mn)            32,709
Excluding cash and Debt            30,377
Value of each share (32,538/486)               62.50
Now lets vary the sales growth from 10% to 7%
Year 2009 2010 2011 2012 2013 2014 2015
Sales (mn)                    19,300            20,651          22,014                  23,379          24,735          26,070          27,374
Growth rate 7.00% 6.60% 6.20% 5.80% 5.40% 5.00%
EBIT (10%)               2,065            2,201                     2,338            2,473            2,607            2,737
Tax (24%)                  496                528                        561                594                626                657
Increase in WC (10% of increae in sales)                  135                136                        136                136                134                130
Free cash flow               1,434            1,537                     1,640            1,744            1,848            1,950
PV of cash flows               1,304            1,270                     1,232            1,191            1,147            1,101
Calculating the current stock price:
Terminal Value of the firm is equal to TV = (FCFn x (1 + g)) / (WACC – g)
in this case we have g as 5% and WACC as 10% and FCF at n as 1950 substituting these in the formula we get =1950*((1+0.05)/(0.1-0.05)) Terminal Value of Nike as: $40950 mn
Now we will compute Present value of all cash flows and terminal value
PV of Terminal Value =40950/(1.1)^6            23,115
Sum of PV of Cash flows above               7,246
Total PV of cash flows (mn)            30,361
Excluding cash and Debt            28,029
Value of each share (32,538/486)               57.67
Now lets vary the sales growth from 10% to 11%
Year 2009 2010 2011 2012 2013 2014 2015
Sales (mn)                    19,300            21,423          23,522                  25,545          27,436          29,137          30,594
Growth rate 11.00% 9.80% 8.60% 7.40% 6.20% 5.00%
EBIT (10%)               2,142            2,352                     2,555            2,744            2,914            3,059
Tax (24%)                  514                565                        613                658                699                734
Increase in WC (10% of increae in sales)                  212                210                        202                189                170                146
Free cash flow               1,416            1,578                     1,739            1,896            2,044            2,179
PV of cash flows               1,287            1,304                     1,307            1,295            1,269            1,230
Calculating the current stock price:
Terminal Value of the firm is equal to TV = (FCFn x (1 + g)) / (WACC – g)
in this case we have g as 5% and WACC as 10% and FCF at n as 2179 substituting these in the formula we get =2179*((1+0.05)/(0.1-0.05)) Terminal Value of Nike as: $45759 mn

Now we will compute Present value of all cash flows and terminal value

PV of Terminal Value =45759/(1.1)^6            25,830
Sum of PV of Cash flows above               7,692
Total PV of cash flows (mn)            33,522
Excluding cash and Debt            31,190
Value of each share (32,538/486)               64.18
For Part B we need to change EBIT margins 7%
Year 2009 2010 2011 2012 2013 2014 2015
Sales (mn)                    19,300            21,230          23,141                  24,992          26,741          28,346          29,763
Growth rate 10% 9% 8% 7% 6% 5%
EBIT (10%)               1,911            2,083                     2,249            2,407            2,551            2,679
Tax (24%)                  459                500                        540                578                612                643
Increase in WC (10% of increae in sales)                  193                191                        185                175                160                142
Free cash flow               1,259            1,392                     1,524            1,654            1,778            1,894
PV of cash flows               1,145            1,150                     1,145            1,130            1,104            1,069
Calculating the current stock price:
Terminal Value of the firm is equal to TV = (FCFn x (1 + g)) / (WACC – g)
in this case we have g as 5% and WACC as 10% and FCF at n as 1894 substituting these in the formula we get =1894*((1+0.05)/(0.1-0.05)) Terminal Value of Nike as: $39775.5 mn

Now we will compute Present value of all cash flows and terminal value

PV of Terminal Value =39775.5/(1.1)^6            22,452
Sum of PV of Cash flows above               6,743
Total PV of cash flows (mn)            29,196
Excluding cash and Debt            26,864
Value of each share (32,538/486)               55.27
Calculating the current stock price:
Terminal Value of the firm is equal to TV = (FCFn x (1 + g)) / (WACC – g)
in this case we have g as 5% and WACC as 10% and FCF at n as 2346 substituting these in the formula we get =2346*((1+0.05)/(0.1-0.05)) Terminal Value of Nike as: $49275.9 mn

Now we will compute Present value of all cash flows and terminal value

PV of Terminal Value =49275.9/(1.1)^6            27,815
Sum of PV of Cash flows above               8,413
Total PV of cash flows (mn)            36,228
Excluding cash and Debt            33,896
Value of each share (32,538/486)               69.75

Attaching screenshots for easier comprehension:-

Cash (mn) Debt (mn) shares o/s (mn) Tax WACC 2300 32 486 24% 10% 2012 24,992 Year Sales (mn) Growth rate EBIT (10%) Tax (24%)

Now lets vary the sales growth from 10% to 7% Year Sales (mn) Growth rate EBIT (10%) Tax (24%) increase in WC (10% of increae

Now lets vary the sales growth from 10% to 11% Year Sales (mn) Growth rate EBIT (10%) Tax (24%) increase in WC (10% of increa

For Part B we need to change EBIT margins 7% Year Sales (mn) Growth rate EBIT (10%) Tax (24%) increase in WC (10% of increae

2012 24,992 2013 Year Sales (mn) Growth rate EBIT (10%) Tax (24%) increase in WC (10% of increae in sales) Free cash flow PV

Thanks